OpenStudy (anonymous):

Two ships were 700 miles apart at midnight and were headed directly toward each other. If they collided at 10 am, find both of the speeds of the ships if one was traveling 30 miles per hour faster than the other.

OpenStudy (anonymous):

Let speed of first =x m/hr speed of second=x+30 m/hr time taken 10 hrs 10x+10(x+30)=700 solve and find the value of x.